Item

8

8

points

eBook

Ask

<span aria-hidden="true" data-icon="p" class="resource__icons" style="margin: 0px; padding: 0px; font: inherit; border: 0px; height: 30px;"></span><span style="margin: 0px; padding: 0px; font: inherit; border: 0px;">Print</span>

Item 8Exercise 5-58

A manufacturer of computer chips claims that the probability of a defective chip is 0.005. The manufacturer sells chips in batches of 700 to major computer companies.

 

a. How many defective chips would you expect in a batch? (Round the final answer to 2 decimal places.)

 

Number of chips            

 

b. What is the probability that none of the chips are defective in a batch? (Round the final answer to 4 decimal places.)

 

Probability            

 

c. What is the probability at least one chip is defective in a batch? (Round the final answer to 4 decimal places.)

 

Probability            




1
Expert's answer
2021-03-09T05:34:55-0500

We have that

p(defective) = 0.005

n = 700

a. Number of chips = n * p = 0.005 * 700 = 3.50

b. P(none of the chips are defective in a batch) = P(X=0)

this follows binomial distribution and can be calculated by the formula:

"P(X=m)=C(n,m)\cdot p^m \cdot (1-p)^{n-m}"

"P(X=0)=C(700,0)\cdot 0.005^0 \cdot (1-0.005)^{700-0}=0.995^{700}=0.0299"

c. P(at least one chip is defective in a batch) = 1 – P(none of the chips are defective in a batch) = 1 – P(X=0) = 1 – 0.0299 = 0.9701


Answer:

a. 3.50

b. 0.0299

c. 0.9701